# An Efficient Difference-of-Convex Solver for Privacy Funnel

### v1

@@ Appendix E Experiment Details
For both the proposed and the compared methods, the deep neural network architectures are described as follows: the encoder is implemented as a multi-layer perceptron (MLP) with $784\rightarrow 500\rightarrow 500\rightarrow(d,d)$ fully-connected (FC) layers where the last pair corresponds to the parameterized mean vector and diagonal covariance matrix of the coded Gaussian vector $\boldsymbol{z}$ similar to the VAE [20]. As for the decoder, our method consists of two parts: 1) the reconstruction block which reverses the architecture of the encoder but starts with $d$ neurons instead. 2) classifier of the coded $\boldsymbol{z}\sim\boldsymbol{\mu}_{x}+\boldsymbol{\varepsilon}\circ\boldsymbol{\sigma}_{x}$ (element-wise), where $\boldsymbol{\varepsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d})$. The linear classifier is implemented as a MLP with $d\rightarrow|Y|$ FC layers, followed by a softmax output activation. As for the compared method, we follow [7] with minimal modification. The same encoder is adopted, but the decoder is modified to a $(|\mathcal{Y}|+d)\rightarrow 500\rightarrow 784$ FC MLP whose inputs are concatenated from $\boldsymbol{z}$ and a one-hot label vector (an elementary vector of $|Y|$ dimensions).
In training the neural networks, we update the weights alternatingly. We fix the weights of the encoder and update the decoder with (17a). Then we fix the weights of the updated decoder and update the encoder parameters (including that of the prior) through (18a). This alternating procedure is implemented in PyTorch with the combination of an on/off requires_grad flag for the parameters of the encoder/decoder and activating the retain_graph flag in the backpropagation of (17a).
@@ Appendix F Extended Evaluation: Compression Ratio and Privacy-Utility Trade-off
Appendix G Extended Evaluation: Visualization
In Fig. 4b, we show the first $64$ samples of the MNIST testing samples with the reconstruction of the proposed DCA solver with $\alpha=0.01$ and $\beta=0.05$. The PSNR for this configuration results is $24$ dB with adversarial success rate of $0.57$. The samples are perceptually recognizable after reconstruction, demonstrating sufficient utility of the images $X$. Then in Fig. 5, we visualize the projected $Z$ of the same configuration. Moreover, we compare the $2$-dimensional projection from our code to direct projection of $X$, i.e., no privatization. As shown from Fig. 5b, projecting our privatized codes has more ambiguity between all classes whereas in Fig. 5a, the direct projection baseline reveals recognizable separation between them instead. This means lower privacy leakage for the proposed approach under sufficient reconstruction quality.

### v3

@@ Appendix E Experiment Details
For both the proposed and compared methods, the deep neural network architectures are described as follows: the encoder is implemented as a multi-layer perceptron (MLP) with $784\rightarrow 500\rightarrow 500\rightarrow(d,d)$ fully-connected (FC) layers where the last pair corresponds to the parameterized mean vector and diagonal covariance matrix of the coded Gaussian vector $\boldsymbol{z}$ similar to the VAE [20]. As for the decoder, our method consists of two parts: 1) the reconstruction block which reverses the architecture of the encoder but starts with $d$ neurons instead. 2) classifier of the coded $\boldsymbol{z}\sim\boldsymbol{\mu}_{x}+\boldsymbol{\varepsilon}\circ\boldsymbol{\sigma}_{x}$ (element-wise), where $\boldsymbol{\varepsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d})$. The linear classifier is implemented as a MLP with $d\rightarrow|Y|$ FC layers, followed by a softmax output activation.
[…]
@@
In training the proposed neural networks prototype, we update the weights alternatingly. We fix the weights of the encoder and update the decoder with (16a). Then we fix the weights of the updated decoder and update the encoder parameters (including that of the prior) through (17a). This alternating procedure is implemented in PyTorch with the combination of an on/off requires_grad flag for the parameters of the encoder/decoder and activating the retain_graph flag in the backpropagation of (16a).
@@ Appendix F Extended Evaluation: Compression Ratio and Privacy-Utility Trade-off
Appendix G Extended Evaluation: Visualization
In Fig. 4, we show the first $64$ samples of the MNIST testing samples with the reconstruction of the proposed DCA solver with $\alpha=0.01$ and $\beta=0.05$. The PSNR for this configuration results is $24$ dB with adversarial success rate of $0.57$. The samples are perceptually recognizable after reconstruction, demonstrating sufficient utility of the images $X$. Then in Fig. 5a and 5b, we visualize the projected $Z$ of the same configuration. We follow the same method described in Section IV-B and use UMAP [28] to project the representation $Z$ to a $\mathbb{R}^{2}$ space with the default hyperparameters adopted. Moreover, we compare the $2$-dimensional projection from our code to direct projection of $X$, i.e., no privatization. As shown from Fig. 5b, projecting our privatized codes has more ambiguity between all classes whereas in Fig. 5a, the direct projection baseline reveals recognizable separation between them instead. This means lower privacy leakage for the proposed approach under sufficient reconstruction quality.